WLCG-OSG-EGEE Operations meeting
→
Europe/Zurich
28-R-15 (CERN conferencing service (joining details below))
28-R-15
CERN conferencing service (joining details below)
Description
grid-operations-meeting@cern.ch
Weekly OSG, EGEE, WLCG infrastructure coordination meeting.
We discuss the weekly running of the production grid infrastructure based on weekly reports from the attendees. The reported issues are discussed, assigned to the relevant teams, followed up and escalated when needed. The meeting is also the forum for the sites to get a summary of the weekly WLCG activities and plans
Attendees:
OSG operations team
EGEE operations team
EGEE ROC managers
WLCG coordination representatives
WLCG Tier-1 representatives
other site representatives (optional)
GGUS representatives
VO representatives ROCs: AsiaPacific, France, Russia, SWEurope,
VOs: CMS, Alice, ALICE, ATLAS
Weekly OSG, EGEE, WLCG infrastructure coordination meeting.
We discuss the weekly running of the production grid infrastructure based on weekly reports from the attendees. The reported issues are discussed, assigned to the relevant teams, followed up and escalated when needed. The meeting is also the forum for the sites to get a summary of the weekly WLCG activities and plans
Attendees:
To dial in to the conference:
a. Dial +41227676000
b. Enter access code 0140768
OR click HERE
NB: Reports were not received in advance of the meeting from:
-
-
16:00
→
16:01
Feedback on last meeting's minutes 1m
-
16:01
→
16:30
EGEE Items 29m
-
<big> Grid-Operator-on-Duty handover </big>From: SouthWestern, UK/Ireland
To: France / CentralEurope
Issues: quiet week, no problems - can't specify anything important as the COD dashboard is inaccessible right now. -
<big> PPS Report & Issues </big>PPS reports were not received from these ROCs:
AP, FR, IT, NE, RU, SEE
Issues from EGEE ROCs:
- SAM UI at Cyfronet has been configured to submit jobs also to uncertified PPS sites(ticket #28436).
There was a problem with permissions on lfc02.pic.es, which cause failure of RM tests for all sites - fixed (ticket #30035). [CE ROC]
Release News:
- gLite 3.1.0 PPS Udate 11 is about to be released to PPS sites (due today)
Notably it contains, together with the usual bug fixes, the new VOBOX service for SL4 (32 bit)
glite-yaim-core has been released, all the metapackages have been updated.
- SAM UI at Cyfronet has been configured to submit jobs also to uncertified PPS sites(ticket #28436).
-
<big> EGEE issues coming from ROC reports </big>
- ROC-DECH
SAM tests for SRM are currently only submitted every two hours. Can this frequency be doubled?
- ROC-DECH
-
<big> gLite Release News</big>gLite 3.1 Update 07 included patch #1389, an update to GFAL/lcg_util. A serious problem has been found with this patch, whereby lcg-cr segfaults with a classic SE endpoint;
Consequently, this patch has been removed from the production repository. Sites which have not yet upgraded will not be affected, but sites which have already upgraded to the affected rpms should do the following
<verbatim> # rpm -e --nodeps GFAL-client lcg_util CGSI_gSOAP_2.7
# yum update glite-WN </verbatim>This will roll back to the earlier versions. For the record, the rpms removed are;
- GFAL-client-1.10.5-1.slc4.i386.rpm
- lcg_util-1.6.4-1.slc4.i386.rpm
- CGSI_gSOAP_2.7-1.2.1-2.i386.rpm
-
<big> Proposal to stop using dteam VO in SAM monitoring </big>For historical reasons some of the standard, regular submissions of SAM tests are being carried out under the dteam VO. We would like to stop doing this and have everything under the OPS VO by next Monday (17 December).
Anyone with an objection to this should contact sam-support@cern.ch before Friday 14 December. -
<big> Upgrade of CERN AFS UI from ggLite 3.0 to gLite 3.1 </big>As announced on Monday, 22 Oct 2007, the default version of the AFS UI will be changed from gLite 3.0 to gLite 3.1.
In practical terms this means that the 'current' AFS UI link will not point to the latest 3.0 version anymore, but to the latest 3.1 version.
The change will happen this Wednesday, 12 Dec, 2007 10:00 CET (09:00 UTC)
An EGEE broadcast will also be sent out today to announce the change.
-
-
16:30
→
17:00
WLCG Items 30m
-
<big> WLCG issues coming from ROC reports </big>
-
<big>WLCG Service Interventions (with dates / times where known) </big>Link to CIC Portal (broadcasts/news), scheduled downtimes (GOCDB) and CERN IT Status Board
Time at WLCG T0 and T1 sites.
-
<big> ATLAS service </big>See also https://twiki.cern.ch/twiki/bin/view/Atlas/TierZero20071 and https://twiki.cern.ch/twiki/bin/view/Atlas/ComputingOperations for more information.
-
<big>CMS service</big>
- Item 1
Speaker: Mr Daniele Bonacorsi (CNAF-INFN BOLOGNA, ITALY) -
<big> LHCb service </big>
- Item 1
Speaker: Dr roberto santinelli (CERN/IT/GD) -
<big> ALICE service </big>
- Item 1
Speaker: Dr Patricia Mendez Lorenzo (CERN IT/GD) -
<big> WLCG Service Coordination </big>
- Item 1
Speaker: Harry Renshall / Jamie Shiers
-
-
17:00
→
17:30
OSG Items 30mSpeaker: Rob Quick (OSG - Indiana University)
-
17:30
→
17:35
Review of action items 5m
-
17:35
→
17:36
AOB 1m
- Item 1
-
16:00
→
16:01